OpenStudy (anonymous):

Simplify the expression. -5+i/2i

OpenStudy (anonymous):

\[\frac{ -5+i }{ 2i }\times \frac{ i }{ i }=\frac{ -5i+i ^{2} }{ 2i ^{2} }=\]

OpenStudy (anonymous):

There's a start. Now use the fact that i^2 = -1 to simplify further.

OpenStudy (anonymous):

what next?